# Estimation of uncertainties in the density driven flow in fractured porous media using MLMC

### v1

@@ 4 Numerical Experiments
We performed numerical experiments with four uncertain parameters: thickness, recharge, permeability and porosity. Permeability and porosity are mutually dependent. To check that uncertainties in all four parameters contribute significantly to the solution, we performed auxiliary experiments with only 1-2 random parameters. First, only one of these four parameters was uncertain, then thickness and recharge were uncertain. In this way we checked that there were no parameters whose contribution could be neglected.
@@
In Figure 2 (top left) we visualise the coefficient of variation $CV(g)$, which is the ratio of the standard deviation to the mean. The variance $V_{\ell}:=\mathbb{V}\left[g_{\ell}\right]$ is shown on the top right, the mean $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ on the bottom left and the variance $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ on the bottom right. The QoI is $g:=c_{m}(t,\mathbf{x}_{1})$ and the time $t=0,\ldots,47\tau$ is along the $x$ axis.
Figure 2: (top left) The coefficient of variance $CV_{\ell}:=CV(g_{\ell})$, (top right) the variance $\mathbb{V}\left[g_{\ell}\right]$, (bottom left) the mean $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$, (bottom right) the variance $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$. The QoI is $g=c_{m}(t,\mathbf{x}_{1})$. The small oscillations in the two lower pictures are due to the dependence of the recharge $\hat{q}_{x}$ on the time, cf. (19).
@@ 4.1 The mean value and variance
Fig. 3 shows the mean value $\mathbb{E}\left[c_{m}(t,\mathbf{x})\right]$ at time points $\{7\tau,19\tau,40\tau,94\tau\}$. In all cases $\mathbb{E}\left[c_{m}(t,\mathbf{x})\right]\in[0,1]$, the dark red color corresponds to the value 1 and dark blue to the value 0.
@@
Figure 3: The mean value $\mathbb{E}\left[c_{m}(t,\mathbf{x})\right]$ at $t=\{7\tau,19\tau,40\tau,94\tau\}$. In all cases, $\mathbb{E}\left[c_{m}\right](t,\mathbf{x})\in[0,1]$.
Figure 4: The variance $\mathbb{V}\left[c_{m}(t,\mathbf{x})\right]$ at $t=\{7\tau,19\tau,40\tau,94\tau\}$. Maximal values (dark red colour) of $\mathbb{V}\left[c_{m}\right]$ are $1.9\cdot 10^{-3}$, $3.4\cdot 10^{-3}$, $2.9\cdot 10^{-3}$, $2.4\cdot 10^{-3}$ respectively. The dark blue colour corresponds to a zero value.
@@ 4.2 Estimation of convergence rates $\alpha$ and $\beta$ for various QoIs
Figure 5 (left) shows the decay of $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ as a function of $g_{\ell}-g_{\ell-1}$. This dependence is fitted by a $2^{-\alpha\ell+\zeta_{1}}$ function. The parameter $\alpha$ indicates the weak convergence rate. Figure 5 (right) shows the decay of $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ with respect to $g_{\ell}-g_{\ell-1}$. This dependence is fitted by a $2^{-\beta\ell+\zeta_{2}}$ curve, where the parameter $\beta$ indicates the strong convergence rate. The constant are $C_{1}=0.47$, $C_{2}=3.7\cdot 10^{-3}$. The QoI is $g:=c_{m}(t_{15},\mathbf{x}_{1})$. The calculated rates $\alpha=1.07$ and $\beta=1.97$ are very close to the theoretical values, which are $1$ and $2$ respectively.
Figure 5: The weak and the strong convergences, QoI is $g:=c_{m}(t_{15},\mathbf{x}_{1})$, $\alpha=1.07$, $\zeta_{1}=-1.1$, $\beta=1.97$, $\zeta_{2}=-8$.
In Fig. 6(left) the decays of $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ and $\mathbb{E}\left[g_{\ell}\right]$ are plotted against the level $\ell$. The QoI is $c_{m}(t_{18},\mathbf{x}_{1})$, $\mathbf{x}_{1}=(1.1,-0.8)$. Similarly, Figure 6 (right) shows that $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ decays much faster than $\mathbb{V}\left[g_{\ell}\right]$.
Figure 6: Decay comparison of (left) $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ and $\mathbb{E}\left[g_{\ell}\right]$ vs. $\ell$; (right) $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ and $\mathbb{V}\left[g_{\ell}\right]$. QoI $g=c_{m}(t_{18},\mathbf{x}_{1})$, $\mathbf{x}_{1}=(1.1,-0.8)$.
In Table 3 we provide estimations for the weak and strong convergence rates $\alpha$ and $\beta$, as well as constants $C_{1}$ and $C_{2}$ for different QoIs. Each QoI is the solution at a point $\mathbf{x}_{i}$ or an integral $I_{i}$ as in (43). We see that the value $\alpha$ is very close to one. This is fully corresponding to the theory. The estimated value of $\beta$ is also very close to the theoretical value, which is equal to $22\alpha=$ 2. The values of $\alpha$ and $C_{1}$ are used for estimation of the number of needed levels $L$, see [31].
@@
Table 3: Estimated the weak and strong convergence rates $\alpha$ and $\beta$, and constants $C_{1}$ and $C_{2}$ for different QoIs. QoI are solution in a point $\mathbf{x}_{i}$ and an integral $I_{i}$ as in (43).
In Table 4 we show the numbers $m_{\ell}$ needed to get the MSE error below $\varepsilon^{2}$, and $\varepsilon$ should be understood as a relative error, as it is defined in (39). All numbers are obtained for the QoI $g=c_{m}(t_{15},\mathbf{x}_{1})$. The numbers $m_{ell}$ can be very different for other QoIs.
@@ 4.3 Comparison of MC and MLMC methods
In Figure 7 we compare the estimated costs of the MC (denoted by a blue dashed line) and MLMC (solid red line) methods, see the two bottom graphics. We also plot the $\varepsilon^{2}$ graphic (yellow dashed line) and the theoretical MC and MLMC costs, purple dashed line and green dashed line respectively. The theoretical MC cost is $\mathcal{O}(\varepsilon^{-2-\frac{\hat{d}\gamma}{\alpha}})$, with $\hat{d}=2+1=3$, $\gamma=1$, $\alpha=1.08$. One can see that from a certain epsilon $\varepsilon\approx 8\cdot 10^{-3}$ the MLMC outperforms the MC method. It can also be seen that the theoretical MC and MLMC lines (two top graphics) have the same slope as the calculated MC and MLMC lines (two bottom graphics). The theoretical MC and MLMC graphics are scaled by a constant for better visibility.
Figure 7: Complexity comparison of ML and MLMC methods against $\varepsilon$ (horizontal axis) in log-log scale.
@@ 5 Discussion and Conclusion
In this work we have considered a setting that mimics the Henry problem [42, 43], modeling seawater intrusion into a 2D coastal aquifer. The pure water recharge from the “land side” resists the salinisation of the aquifer due to the influx of saline water through the “sea side”, thereby achieving some equilibrium in the salt concentration. In our setting, following [20], we consider a fracture on the sea side that significantly increases the permeability of the porous medium.
@@
The flow and transport essentially depend on the geological parameters of the porous medium, including the fracture. We investigated the effects of various uncertainties on saltwater intrusion. We assumed uncertainties in the fracture width, the porosity of the bulk medium, its permeability and the pure water recharge from the land side. The porosity and permeability were modeled by random fields, the recharge by a random but periodic intensity and the thickness by a random variable. We calculated the mean and variance of the salt mass fraction, which is also uncertain.
The main question we investigated in this work was how well the MLMC method can be used to compute statistics of different QoIs. We found that the answer depends on the choice of the QoI. First, not every QoI requires a hierarchy of meshes and MLMC. Second, MLMC requires stable convergence rates for $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ and $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$. These rates should be independent of $\ell$. If these convergence rates vary for different $\ell$, then it will be hard to estimate $L$ and $m_{\ell}$, and MLMC will either not work or be suboptimal. We were not able to get stable convergence rates for all levels $\ell=1,\ldots,5$ when the QoI was an integral as in (43). We found that for $\ell=1,\ldots 4$ and $\ell=5$ the rate $\alpha$ was different. Further investigation is needed to find the reason for this. Another difficulty is the dependence on time, i.e. the number of levels $L$ and the number of sums $m_{\ell}$ depend on $t$. At the beginning the variability is small, then it increases, and after the process of mixing salt and fresh water has stopped, the variance decreases again.
The number of random samples required at each level was estimated by calculating the decay of the variances and the computational cost for each level. These estimates depend on the minimisation function in the MLMC algorithm.
@@
To achieve the efficiency of the MLMC approach presented in this work, it is essential that the complexity of the numerical solution of each random realisation is proportional to the number of grid vertices on the grid levels. This holds only under special choice the numerical solvers, and we applied the geometric multigrid method implemented in the ug4 software toolkit, cf. [37, 49]. Our numerical tests confirm the expected efficiency. The random realizations do not depend on each other and have been computed concurrently on different nodes of the supercomputer Shaheen II at the King Abdullah University of Science and Technology.
In the numerical experiments, we computed the expected value and variance of the mass fraction in the whole domain and the solution at a few preselected points $(t,\mathbf{x})$. It turns out that the grid hierarchy of only 2 to 3 levels is sufficient for some QoIs. Note that a different type of porosity in (20) may lead to a different conclusion.
Our numerical results confirm that the MLMC is principally more numerically efficient than the MC method. Thus, sampling at different mesh levels helps to reduce the overall computational cost.
@@
Constraints. 1. The time dependence is challenging. The optimal number of samples depends on the time moment $t$, and the spatial location $\mathbf{x}$, and may be small for some points and large for others. 2. Each new QoI requires new estimates of all parameters $\alpha$, $\beta$, $\gamma$, and the resulting MLMC performance graphics can be different.
Future work. A more realistic modeling of the fracture, porosity and permeability with more random variables can be considered.

### v3

@@ 4 Numerical Experiments
The simulations below depend on three stochastic variables $\{\xi_{1},\xi_{2},\xi_{3}\}$, $\xi_{1}$ is controlling thickness of the fracture (18), $\xi_{2}$ the values of permeability and porosity (20), and $\xi_{3}$ the intensity of the recharge, as described in (19).
Thus, in total there are four uncertain parameters: thickness, recharge, permeability and porosity. Permeability and porosity are mutually dependent. To verify that the uncertainties in all four parameters contribute significantly to the solution, we performed auxiliary experiments. These experiments showed that all parameters contribute significantly and that there are no parameters whose contribution could be neglected.
@@
In Figure 2 (top left) we visualise the coefficient of variation $CV_{\ell}(g):=\frac{\sigma(g_{\ell})}{\mathbb{E}\left[g_{\ell}\right]}$, which is the ratio of the standard deviation $\sigma(g_{\ell})$ to the mean $\mathbb{E}\left[g_{\ell}\right]$. The variance $V_{\ell}:=\mathbb{V}\left[g_{\ell}\right]$ is shown on the top right. The mean value of the difference $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ is shown on the bottom left and the variance $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ on the bottom right. The QoI is $g:=c_{m}(t,\mathbf{x}_{1})$ and the time $t=0,\ldots,47\tau$ is along the $x$ axis.
Figure 2: (top left) The coefficient of variance $CV_{\ell}:=CV(g_{\ell})$, (top right) the variance $\mathbb{V}\left[g_{\ell}\right]$, (bottom left) the mean $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$, (bottom right) the variance $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$. The QoI is $g=c_{m}(t,\mathbf{x}_{1})$. The small oscillations in the two lower pictures are due to the dependence of the recharge $\hat{q}_{\mathrm{in}}$ on the time, cf. (19). The time is along the $x$ axis, $t\in[0,47\tau]$
@@ 4.1 The mean value and variance
Figure 3 shows the mean value $\mathbb{E}\left[c_{m}(t,\mathbf{x})\right]$ at time points $\{7\tau,19\tau,40\tau,94\tau\}$. In all cases $\mathbb{E}\left[c_{m}(t,\mathbf{x})\right]\in[0,1]$, the dark red color corresponds to the value 1 and dark blue to the value 0. The fracture is visible in the center, too. Due to the high permeability inside the fracture for the most realizations, the flow is divided into two subdomains (compare the red areas above and below the fracture). The high-permeable fracture isolates the upper flow from the lower flow. If we further increase the permeability inside the fracture, we get two almost independent problems: one above the fracture and one below.
@@
Figure 3: The mean value $\mathbb{E}\left[c_{m}(t,\mathbf{x})\right]$ at $t=\{7\tau,19\tau,40\tau,94\tau\}$. In all cases, $\mathbb{E}\left[c_{m}\right](t,\mathbf{x})\in[0,1]$. The blue colour corresponds to $\mathbb{E}\left[c_{m}\right](t,\mathbf{x})=0$ and the dark red colour to $\mathbb{E}\left[c_{m}\right](t,\mathbf{x})=1$.
Figure 4: The variance $\mathbb{V}\left[c_{m}(t,\mathbf{x})\right]$ at $t=\{7\tau,19\tau,40\tau,94\tau\}$. Maximal values (dark red colour) of $\mathbb{V}\left[c_{m}\right]$ are $1.9\cdot 10^{-3}$, $3.4\cdot 10^{-3}$, $2.9\cdot 10^{-3}$, $2.4\cdot 10^{-3}$ respectively. The blue colour corresponds to a zero value.
@@ 4.2 Estimation of convergence rates $\alpha$ and $\beta$ for various QoIs
Figure 5 (left) shows the decay of $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ as a function of $g_{\ell}-g_{\ell-1}$. The $0y$ axis denotes $\log_{2}(\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right])$. This dependence is fitted by a $2^{-\alpha\ell+\zeta_{1}}$ function. The parameter $\alpha$ indicates the weak convergence rate. Figure 5 (right) shows the decay of $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ with respect to $g_{\ell}-g_{\ell-1}$. The $0y$ axis denotes $\log_{2}(\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right])$. This dependence is fitted by a $2^{-\beta\ell+\zeta_{2}}$ curve, where the parameter $\beta$ indicates the strong convergence rate. The constants are $C_{1}=0.47$ and $C_{2}=3.7\cdot 10^{-3}$. The QoI is $g:=c_{m}(t_{15},\mathbf{x}_{1})$. The calculated rates $\alpha=1.07$ and $\beta=1.97$ are very close to the theoretical values, which are $1$ and $2$ respectively.
Figure 5: The weak and the strong convergences, QoI is $g:=c_{m}(t_{15},\mathbf{x}_{1})$, $\alpha=1.07$, $\zeta_{1}=-1.1$, $\beta=1.97$, $\zeta_{2}=-8$. The $0y$ axis denotes $\log_{2}(\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right])$ on the left and $\log_{2}(\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right])$ on the right.
In Fig. 6 (left) the decays of $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ and $\mathbb{E}\left[g_{\ell}\right]$ are plotted against the level $\ell$. The QoI is $c_{m}(t_{15},\mathbf{x}_{1})$, $\mathbf{x}_{1}=(1.1,-0.8)$. The values on the $0y$ axis are $\log_{2}(\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right])$ and $\log_{2}(\mathbb{E}\left[g_{\ell}\right])$. Similarly, Figure 6 (right) shows that $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ decays much faster than $\mathbb{V}\left[g_{\ell}\right]$. The values on the $0y$ axis are $\log_{2}(\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right])$ and $\log_{2}(\mathbb{V}\left[g_{\ell}\right])$.
Figure 6: Decay comparison of (left) $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ and $\mathbb{E}\left[g_{\ell}\right]$ vs. $\ell$; (right) $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ and $\mathbb{V}\left[g_{\ell}\right]$. QoI $g=c_{m}(t_{15},\mathbf{x}_{1})$, $\mathbf{x}_{1}=(1.1,-0.8)$. The $0y$ axis shows $\log_{2}(\cdot)$ of $\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$, $\mathbb{E}\left[g_{\ell}\right]$, $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$, and $\mathbb{V}\left[g_{\ell}\right]$ values.
In Table 3 we provide estimations for the weak and strong convergence rates $\alpha$ and $\beta$, as well as constants $C_{1}$ and $C_{2}$ for different QoIs. Each QoI is the solution at a point $\mathbf{x}_{i}$ or an integral $I_{i}$ as in (42). We see that the value $\alpha$ for $c(\mathbf{x}_{i},15\tau)$ is very close to $-1.0$, and $-1.5$ for $I_{2}$. The value $\alpha=1.0$ is fully corresponding to the theory. The estimated value of $\beta$ is also very close to the theoretical value, which is $2\alpha\approx 2$. The values of $\alpha$ and $C_{1}$ are used for estimation of the number of needed levels $L$, see also [35]. The value $\alpha=-1.5$ for $I_{2}$ is also what we expected, since the integral of $c$ is usually a smoother function than just $c$. The value $\beta=-2.6$ (in the last row) is also close to the theoretical value, which is $32\alpha=-$ 3.
@@
Table 3: Estimated the weak and strong convergence rates $\alpha$ and $\beta$, and constants $C_{1}$ and $C_{2}$ for different QoIs. QoI are solution in a point $\mathbf{x}_{i}$ and an integral $I_{2}$ as in (42).
In Table 4 we show the numbers $m_{\ell}$ needed to get the MSE error below $\varepsilon^{2}$, and $\varepsilon$ should be understood as a relative error, as it is defined in (38). All numbers are obtained for the QoI $g=c_{m}(t_{15},\mathbf{x}_{1})$. The numbers $m_{\ell}$ can be very different for other QoIs.
@@ 4.3 Comparison of MC and MLMC methods
In Figure 7(left) we compare the estimated costs of the MC (denoted by a blue solid line) and MLMC (red solid line) methods. The cost of MLMC is calculated as in (34). We only give the estimated cost of MC because it is too time-consuming to compute. Such a comparison is typical for MLMC papers [10, 25, 26]. In addition, in Figure 7(right) we plot the slopes of the $\varepsilon^{2}$ graph (yellow solid line) and the theoretical MC and MLMC costs, purple dotted line and red dashed line, respectively. Note that these three graphs are not scaled to make them easier to read, only the slope is important. The theoretical MC cost is $\mathcal{O}(\varepsilon^{-2-\frac{\hat{d}\gamma}{\alpha}})$, with $\hat{d}=2+1=3$, $\gamma=1$ (linear cost of the geometric multigrid method), $\alpha=1.08$. Thus, the theoretical MC cost is $\mathcal{O}(\varepsilon^{-5})$. The cost of MLMC, according to the third case in (36), where $\beta<\tilde{d}\cdot\gamma$, is $\mathcal{O}(\varepsilon^{-(2+\frac{3\cdot 1-2}{1})})=\mathcal{O}(\varepsilon^{-3})$. This graph is also up to a constant factor. From the left figure, we can see that the MLMC outperforms the MC method for almost all values of $\varepsilon$. The right figure also shows that the theoretical MC and MLMC lines (two top plots) have the same slope as the computed MC and MLMC lines (two bottom plots).
Figure 7: (left) Computational cost comparison of MC and MLMC methods against $\varepsilon$ (horizontal axis) in log-log scale. (right) Same as on the left with additional graphs showing the slopes of the theoretical costs. The cost (in seconds) is on the $y$ axis.
@@ 5 Discussion and Conclusion
In this study we examined a setting that mimics the Henry-like problem [45, 46], modeling seawater intrusion into a 2D coastal aquifer. The solution shows how pure water recharge (see Fig. 1(left)) from the “land side” mitigates the salinization of the aquifer due to the influx of saline water from the “sea side,” thereby achieving an equilibrium in salt concentration. Building on [21], we incorporated a fracture on the seaside that significantly enhances the permeability of the porous medium.
@@
Uncertainties were introduced in the fracture width, the porosity of the bulk medium, its permeability and the pure water recharge from the land side. Porosity and permeability were modelled as random fields, recharge as a random periodic intensity and fracture width as a random variable. In total, three random variables were involved. We calculated the mean and variance of the salt mass fraction, which is subject to uncertainty.
The central question of this study was to evaluate the efficiency of the Multi-Level Monte Carlo (MLMC) method in computing statistics for different quantities of interest (QoIs). Our results indicate that the efficiency of MLMC depends on the QoI $g_{\ell}$, on the convergence rates of $\mathbb{E}[g_{\ell}-g_{\ell-1}]$ and $\mathbb{V}\left[g_{\ell}-g_{\ell-1}\right]$ (see Table 3 and Fig. 6), which should remain consistent across different levels $\ell$. In our numerical tests, which we do not publish here, we found that for some QoIs the convergence rates change, i.e. do not remain consistent across levels $\ell=1,\ldots,5$. One such example is the integral $I_{4}(c_{m})$, as given in (42). In particular, the rate $\alpha$ for levels $\ell=1,\ldots,4$ was equal to $-1.5$, but for $\ell=5$ it was a completely different number. A possible explanation is that the 5th level is not needed. This observation needs further investigation.
Another challenge was the time dependence. The number of levels $L$ and the number of samples $m_{\ell}$ depend on the decay of the mean and variance (see (40)-(41) and Figures 2 bottom left and bottom right), which in turn depend on time $t$. Initially, when $t$ was small, the variability was low, increased during the mixing process of salt and fresh water, and decreased when equilibrium was reached. So, it is impossible to provide $L$ and $m_{\ell}$, which are optimal for every time step.
The number of samples required at each level was estimated by minimising the objective function (see (30)) used in the MLMC algorithm. A different objective function has been used in some other MLMC work (see [10]). The choice of the goal function can affect the MLMC efficiency.
@@
To achieve the efficiency of the MLMC approach demonstrated in this study, it is crucial that the complexity of solving each random realization scales with the number of grid vertices at each level. This was achieved using the geometric multigrid method implemented in the ug4 software toolkit [41, 50], and our numerical tests validated the expected efficiency. The random realizations were computed concurrently on different nodes of the Shaheen II supercomputer at the King Abdullah University of Science and Technology.
Our results confirm that the MLMC method is applicable to the problem at hand, computing the mean up to 100 times faster than standard Monte Carlo methods (see Fig. 7, left). Sampling across different mesh levels effectively reduces the overall computational cost.
@@
Constraints. Time dependence poses challenges, as the optimal number of samples varies with time $t$ and spatial location $\mathbf{x}$, potentially requiring fewer samples at some points and more at others. Each new QoI necessitates re-evaluation of parameters $\alpha$ and $\beta$, resulting in different MLMC performance characteristics for each QoI.
Future Work. Further research could focus on more realistic modeling of fractures, porosity, and permeability by incorporating additional random variables. One could allow more fractures, their uncertain orientations. One can further research how geological parameters, including the fracture, influence flow and transport.